# Raw2Drive: Reinforcement Learning with Aligned World Models for End-to-End Autonomous Driving

### v1

@@ 1 Introduction
footnotetext: Correspondence author is also affiliated with Shanghai lnnovation Institute. This work was in part supported by by NSFC (62222607) and Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102.
@@
Beyond modular systems, end-to-end autonomous driving (E2E-AD) models chen2024end ; Chitta2023PAMI ; wu2023policypretrainingautonomousdriving ; li2023delving are emerging, where a unified model directly uses raw sensor inputs for planning. As shown in Figure 1, most of these models hu2023planning ; jiang2023vad ; jia2022multi ; jia2023towards are based on imitation learning (IL), which trains models to mimic expert demonstrations. However, IL faces fundamental limitations such as poor generalization to unseen situations wen2020fighting ; zhai2023ADMLP ; wu2022trajectoryguided ; jia2021ide and causal confusion d2014confusion ; jia2023thinktwice ; jia2023driveadapter , which occur when the model incorrectly associates actions with the wrong causes. These issues become particularly problematic in complex and dynamic driving scenarios where decision-making must account for delicate interactions with environment.
Reinforcement learning (RL) mnih2015human offers a promising alternative by optimizing driving policies through reward-driven interactions with the dynamic environment schulman2017proximal ; mnih2013playing , beyond the expert demonstration in imitation learning. Recently, RL has been shown to achieve higher performance bounds than IL in various domains, such as AlphaZero silver2017mastering ; niu2024lightzero , and OpenAI-O3/DeepSeek-R1 guo2025deepseek . These successes highlight RL’s ability to adapt intelligent policies for complex decision-making tasks.
@@
 We only use 64 H800 GPU days in total to deliver our final planner, and the cost can be further saved to 40 GPU days when Think2Drive is reused which dismisses our phase I training. In comparison, IL-based UniAD costs about 30 GPU days yet it only solves even 3$\sim$4 corner cases in CARLA v2. We believe it is significantly less than the imitation learning approaches used in industry [^1] As the companies have never officially reported their technical details, there are only unverified information that Tesla Full-scale Driving (FSD), and Huawei Advanced Driving System (ADS) were both trained by imitation learning by leveraging their massive human driving logs. and the hope is our work could provide an orthogonal reference for industry applications.
@@ 3 Method
Table 2: Notations of Dual Stream World Models in Raw2Drive. WM means the privileged world model and $\hat{\text{WM}}$ means the raw sensor world model. $t$ denotes the time-step. Both world models consist of an Encoder, RSSM, and Heads. The privileged world model is similar to DreamerV3 hafner2023mastering while raw sensor world model has a tailored encoder $\hat{Enc}$ and only has a tailored decoder head $\hat{Dec}$. circled-1 and circled-2 represent the operation during training and inference respectively.
Dual Stream Privileged WM Raw Sensor ${\hat{\text{WM}}}$
@@
World Model RSSM Deterministic State $h_{t}=f_{\theta}(h_{t-1},a_{t-1},s_{t-1})$
@@
The rationale behind the two stream design is that: raw sensor data (e.g., high resolution multi-view outdoor videos with motion) are high-dimensional and complex, presenting significant challenges to directly train a world model with low error brooks2024video ; xia2024rgbd . Thus, rather than building the raw sensor end-to-end world model from scratch, we first build the privileged stream as an auxiliary since the learning of the world model under structured and low-dimensional conditions is much easier and there are already some successes li2024think . The raw sensor stream’s learning process is guided by the trained privileged stream, which eases the task of extracting the decision-making related information from raw sensor data. The Guidance Mechanism is carefully designed to avoid cumulative error and train-val gaps.
@@ 3.3 Guidance Mechanism
In the previous section, we introduce Raw2Drive’s dual-stream MBRL framework. In this section, we give details of the guidance mechanism about (I) how the privileged stream guide the learning process of both the raw sensor world and policy; (II) how to alleviate cumulative errors and train-val gaps of the dual-stream paradigm.
@@
Rollout Guidance. In MBRL framework, the rollout of world model is crucial for effective policy learning. Thus, under the proposed dual-stream framework, the fisrt type of guidance is Rollout Guidance, which adopts the trained privileged world model provides supervisions during the entire rollout process of the raw sensor world model. Specifically, as shown in Figure 6, there are three components in the rollout process: (1) Encoded State $\hat{e}$, (2) Deterministic State $\hat{h}$, (3) Stochastic State $\hat{s}$. At each timestep $t$, both stream encode the observation (either privileged or raw sensor) into Encoded State. However, due to the high dimensionality and inherent redundancy of multi-view video data, its encoded state $\hat{e}$ of the raw sensor stream may exhibit instability and errors. To address this, we introduce a loss function to align it with the privileged encoded state $e$, ensuring spatial consistency at each timestep, called Spatial-Temporal Alignment Loss.
@@ 4.1 Datasets and Benchmark
Table 3: Performance on Carla Official Town13 Validation and Devtest Benchmark. *denotes expert feature distillation. The online Leaderboard 2.0 closed since October 2024. As discussed in carla-garage Zimmerlin2024ArXiv and Section 4.3, long routes evaluation in Leaderboard 2.0 can’t reflect the actual driving performance jia2024bench .
@@ 4.2 Metric
DS $\uparrow$ SR(%) $\uparrow$ Efficiency $\uparrow$ Comfort $\uparrow$
@@ 4.4 Ablation Study
Choice of Heads in Raw Sensor World Model.  Due to the high-dimensional redundancy of image information and the fact that both the reward and continue flags are represented as scalars, the network struggles to achieve stable convergence during training. To further investigate this, we conduct an ablation study on the raw sensor world model heads. As shown in Table 7, the decoder head provides effective supervision, significantly enhancing the learning of the raw sensor world model. In contrast, the reward and continuation heads introduce ambiguity that hinders convergence, leading to suboptimal world model training and degraded policy performance. We provide a detailed visualization in the Appendix G.2.
@@
Table 7: Ablation on the Three Latent State Alignments in Rollout Guidance.
@@
Rollout Guidance. We conduct an ablation study to examine the role of rollout guidance in the learning process of the raw sensor world model. As shown in Table 7, since the world model is trained during rollouts, any state misalignment can lead to discrepancies between the predictions of the two world models. Our results show that when any component of the rollout guidance is missing, the model can only handle simple tasks, such as moving straight or turning left, while failing to address more complex interactive corner cases. This misalignment significantly hinders the training of the raw sensor-based world model. All three components of rollout guidance components are thus essential for robust learning and performance. Loss comparisons are provided in Appendix F.
@@ 5 Conclusion
We propose Raw2Drive, in public literature, the first end-to-end model-based reinforcement learning method in autonomous driving. Our approach introduces a novel dual-stream architecture and a guidance mechanism to effectively enable MBRL learning. The approach is fulfilled by the careful treatment of the two world models for the raw and privileged information, respectively, and achieve the state-of-the-art performance on lately released benchmarks. Furthermore, our proposed paradigm is not limited to camera-only systems and is applicable to guiding the learning of any sensor through the privileged input. We hope this work serves as a stepping stone toward exploring reinforcement learning for end-to-end autonomous driving.
@@ A.1 End-to-End Autonomous Driving Based on Imitation Learning
Imitation Learning (IL) chen2024end ; pan2017agile ; jia2024amp ; yang2025trajectory ; zhu2024flatfusion ; lu2024activead ; yang2023llm4drive is a foundational method for training agents by mimicking expert behavior, especially suitable for end-to-end autonomous driving tasks. IL aims to replicate human driving actions, thus reducing system complexity and simplifying the decision-making process from perception to control.
@@ A.2 End-to-End Autonomous Driving based on Model-free Reinforcement Learning
Model-free reinforcement learning zhang2021end ; toromanoff2020end is another key approach for end-to-end autonomous driving, optimizing decision-making through rewards rather than direct imitation. Early RL work, such as MaRLn toromanoff2020end , demonstrated the potential of reinforcement learning in autonomous driving. However, it required extensive pre-training, with around 20M (23 days) of training in a single town or 50M (57 days) in multi-towns to solve only 4 standard cases in CoRL 2017 dosovitskiy2017carla . The complexity of perception in autonomous driving makes it difficult to optimize perception and control jointly, increasing the challenge of training end-to-end systems with RL. To reduce the impact of perception on decision training, some classic RL methods, such as Roach zhang2021end , use 2D Bird’s Eye View (BEV) representations as input observations, simplifying the task of learning from raw sensor data and focusing on high-level environmental features. And then use RL methods to collect offline data for training the end-to-end method wu2022trajectoryguided ; shao2023safety ; shao2023reasonnet .
@@ Appendix B Details of Training Pipeline
Privileged Replay Buffer $B$, Replay Buffer $\hat{B}$ Training iterations $N$
@@ Appendix F Rollout Guidance
Appendix G Visualization
G.1 Decoder Output by the World Model
@@ G.2 Perception Confusion
G.3 Limitations and Social Impact
Limitations: This is the first attempt at model-based reinforcement Learning in end-to-end autonomous driving. In our setting, the privileged input is ground truth bounding boxes and HD-Map. And in real-world autonomous driving, for industry, the ground truth bounding boxes and HD-Map can be from human annotation or advanced perception algorithms. Reinforcement learning in the real world is also a technical issue that would be solved by 3DGS or a diffusion-based simulator in the future. In our work, we explore the world model training guidance from the privileged stream to the raw sensor stream. And we will explore the behavior policy training guidance from privileged stream to raw sensor stream in future work.
Social Impact: Raw2Drive introduces an efficient and effective reinforcement learning framework for end-to-end autonomous driving, addressing critical limitations of imitation learning, such as causal confusion and distribution shift. By enabling the autonomous system to learn robust world models from raw sensor inputs and make informed decisions through reinforcement learning, Raw2Drive’s paradigm has the potential to significantly enhance the safety, reliability, and generalization capabilities of autonomous driving systems.
G.4 Driving Performance Demonstration by Video
See the supplementary material Raw2Drive.mov file for details.

### v2

@@ 1 Introduction
footnotetext: Correspondence Author. This work was in part supported by by NSFC (62222607,62506229) and Natural Science Foundation of Shanghai (25ZR1402268).
@@
Beyond modular systems, end-to-end autonomous driving models chen2024end ; Chitta2023PAMI ; wu2023policypretrainingautonomousdriving ; li2023delving are emerging, where a unified model directly uses raw sensor inputs for planning. As shown in Figure 1, most of these models hu2023planning ; jiang2023vad ; jia2022multi ; jia2023towards are based on imitation learning (IL), which trains models to mimic expert demonstrations. However, IL faces fundamental limitations such as poor generalization to unseen situations wen2020fighting ; zhai2023ADMLP ; wu2022trajectoryguided ; jia2021ide and causal confusion d2014confusion ; jia2023thinktwice ; jia2023driveadapter , which occur when the model incorrectly associates actions with the wrong causes. These issues become problematic in complex and dynamic driving scenarios where decision-making must account for delicate interactions with environment.
Reinforcement learning (RL) mnih2015human ; niu2023lightzero ; pu2024unizero offers a promising alternative by optimizing driving policies through reward-driven interactions with the dynamic environment schulman2017proximal ; mnih2013playing ; Bai_Liu_Du_Wen_Yang_2025 , beyond the expert demonstration in imitation learning. Recently, RL has been shown to achieve higher performance bounds than IL in various domains, such as AlphaZero silver2017mastering ; niu2024lightzero , and OpenAI-O3/DeepSeek-R1 guo2025deepseek . These successes highlight RL’s ability to adapt policies for complex decision-making tasks.
@@
[…]
@@ 3 Method
Table 2: Notations of Dual Stream World Models in Raw2Drive. WM means the privileged world model and $\hat{\text{WM}}$ means the raw sensor world model. $t$ denotes the time-step. Both world models consist of an Encoder, RSSM, and Heads. The privileged world model is similar to DreamerV3 hafner2023mastering while raw sensor world model has a tailored encoder $\hat{Enc}$ and only has a tailored decoder head $\hat{Dec}$. and represent the operation during training and inference respectively.
Dual Stream Privileged WM Raw Sensor $^{\hat{\text{WM}}}$
@@
World
Model
RSSM Deterministic State $h_{t}=f_{\theta}(h_{t-1},a_{t-1},s_{t-1})$
@@
The rationale behind the two stream design is that: raw sensor data (e.g., high resolution multi-view videos) are high-dimensional and complex, presenting significant challenges to directly train a world model with low error brooks2024video ; xia2024rgbd . Thus, rather than building the raw sensor world model from scratch, we first build the privileged stream as an auxiliary since the learning of the world model under structured and low-dimensional conditions is much easier and there are already some successes li2024think . The raw sensor stream’s learning process is guided by the trained privileged stream, which eases the task of extracting the decision-making related information from raw sensor data. The Guidance Mechanism is carefully designed to avoid cumulative error and train-val gaps.
@@ 3.3 Guidance Mechanism
In the previous section, we introduced Raw2Drive’s dual-stream MBRL framework. In this section, we give details of the guidance mechanism about (I) how the privileged stream guide the learning process of both the raw sensor world and policy; (II) how to alleviate cumulative errors and train-val gaps of the dual-stream paradigm.
@@
Rollout Guidance. In MBRL framework, the rollout of world model is crucial for effective policy learning. Thus, under the proposed dual-stream framework, the first type of guidance is Rollout Guidance, which adopts the trained privileged world model provides supervisions during the entire rollout process of the raw sensor world model. Specifically, as shown in Figure 6, there are three components in the rollout process: (1) Encoded State $\hat{e}$, (2) Deterministic State $\hat{h}$, (3) Stochastic State $\hat{s}$. At each timestep $t$, both stream encode the observation (either privileged or raw sensor) into Encoded State. However, due to the high dimensionality and inherent redundancy of multi-view video data, its encoded state $\hat{e}$ of the raw sensor stream may exhibit instability and errors. To address this, we introduce a loss function to align it with the privileged encoded state $e$, ensuring spatial consistency at each timestep, called Spatial-Temporal Alignment Loss.
@@ 4.1 Datasets and Benchmark
Table 3: Performance on Carla Official Town13 Validation and Devtest Benchmark. *denotes expert feature distillation. As discussed in carla-garage Zimmerlin2024ArXiv and Section 4.3, long routes evaluation in Leaderboard 2.0 can’t reflect the actual driving performance jia2024bench .
@@ 4.2 Metric
DS $\uparrow$
SR(%) $\uparrow$
Efficiency $\uparrow$
Comfort $\uparrow$
@@ 4.4 Ablation Study
Choice of Heads in Raw Sensor World Model.  Due to the high-dimensional redundancy of image information and the fact that both the reward and continue flags are represented as scalars, the network struggles to achieve stable convergence during training. To further investigate this, we conduct an ablation study on the raw sensor world model heads. As shown in Table 7, the decoder head provides effective supervision, significantly enhancing the learning of the raw sensor world model. In contrast, the reward and continuation heads introduce ambiguity that hinders convergence, leading to suboptimal world model training and degraded policy performance. We provide a detailed visualization in the Appendix H.2.
@@
Table 7: Ablation on the Abstract-State Alignment in Rollout Guidance.
@@
Abstract-State Alignment in Rollout Guidance. We conduct an ablation study to examine the role of Abstract-State Alignment in rollout guidance. As shown in Table 7, since the world model is trained during rollouts, any state misalignment can lead to discrepancies between the predictions of the two world models. Our results show that when any component of the rollout guidance is missing, the model can only handle simple tasks, such as moving straight or turning left, while failing to address more complex interactive corner cases. This misalignment significantly hinders the training of the raw sensor-based world model. All three components of rollout guidance components are thus essential for robust learning and performance. Loss comparisons are provided in Appendix F.
Spatial-Temporal Alignment in Rollout Guidance. The results on Spatial-Temporal Alignment (based on Dev10) is shown in Table 8. Spatial Alignment ensures consistency between image representations and BEV representations—removing it is similar to “driving blind.” Temporal Alignment maintains the consistency of future predictions across time. Both components are essential and complementary for a stable world model rollout.
Table 8: Ablation on the Spatial-Temporal Alignment in Rollout Guidance.
[…]
$\times$ $\times$ 0.0 0.0/10
$\checkmark$ $\times$ 13.6 1.2/10
$\times$ $\checkmark$ 9.24 0.8/10
$\checkmark$ $\checkmark$ 83.5 7.5/10
@@
Real-time Inference.  We conducted latency analysis for each module. In end-to-end autonomous driving, the perception backbone (e.g., surround-view image encoder) typically dominates the overall latency. Our world model and policy are highly efficient (each under 2ms), while the raw sensor stream is mainly bottlenecked by the vision encoder (e.g., BEVFormer). The results in Appendix G.
@@ 5 Conclusion
We propose Raw2Drive, the first end-to-end model-based reinforcement learning method in autonomous driving. Our approach introduces a novel dual-stream architecture and a guidance mechanism to effectively enable MBRL learning. The approach is fulfilled by the careful treatment of the two world models for the raw and privileged information, respectively, and achieve the state-of-the-art performance on lately released benchmarks. We hope this work serves as a stepping stone toward exploring reinforcement learning for end-to-end autonomous driving.
Limitations: In our setting, the privileged input is ground truth bounding boxes and HD-Map. And in real-world autonomous driving, for industry, the ground truth bounding boxes and HD-Map can be from human annotation or advanced perception algorithms. Reinforcement learning in the real world is also a technical issue that would be solved by 3DGS Ma2025BezierGS ; huang2023iddr ; chen2025focused or a diffusion-based simulator yang2025resim in the future. While CARLA remains the only viable closed-loop simulator for RL research at present, our work focuses on policy learning and introduces a dual-stream world model design that is conceptually decoupled from the specific simulator.
Social Impact: Raw2Drive presents an efficient reinforcement learning framework for end-to-end autonomous driving, mitigating key issues of imitation learning such as causal confusion and distribution shift. By learning robust world models from raw sensors, it enhances the safety, reliability, and generalization of autonomous driving systems.
@@ A.1 End-to-End Autonomous Driving Based on Imitation Learning
Imitation Learning (IL) chen2024end ; pan2017agile ; jia2024amp ; yang2025trajectory ; zhu2024flatfusion ; lu2024activead ; yang2023llm4drive ; yang2025drivemoe is a foundational method for training agents by mimicking expert behavior, especially suitable for end-to-end autonomous driving tasks. IL aims to replicate human driving actions, thus reducing system complexity and simplifying the decision-making process from perception to control.
@@ A.2 End-to-End Autonomous Driving based on Model-free Reinforcement Learning
Model-free reinforcement learning zhang2021end ; toromanoff2020end ; NEURIPS2022_8be9c134 ; Bai_Zhang_Tao_Wu_Wang_Xu_2023 ; bai2024efficient ; bai2025retrieval ; zhang2025amulet ; li2023normalization ; zhang2023gobigger ; zhu2025flowrl is another key approach for end-to-end autonomous driving, optimizing decision-making through rewards rather than direct imitation. Early RL work, such as MaRLn toromanoff2020end , demonstrated the potential of reinforcement learning in autonomous driving. However, it required extensive pre-training, with around 20M (23 days) of training in a single town or 50M (57 days) in multi-towns to solve only 4 standard cases in CoRL 2017 dosovitskiy2017carla . The complexity of perception in autonomous driving makes it difficult to optimize perception and control jointly, increasing the challenge of training end-to-end systems with RL. To reduce the impact of perception on decision training, some classic RL methods, such as Roach zhang2021end , use 2D Bird’s Eye View (BEV) representations as input observations, simplifying the task of learning from raw sensor data and focusing on high-level environmental features. And then use RL methods to collect offline data for training the end-to-end method wu2022trajectoryguided ; shao2023safety ; shao2023reasonnet .
@@ Appendix B Details of Training Pipeline
Privileged Replay Buffer $B$, Replay Buffer $\hat{B}$ Training iterations NN
@@ Appendix F Rollout Guidance
Appendix G Inference Latency
Table 13: Latency Comparison between Privileged and Raw Sensor Streams.
Method Modality Encoder Latency (ms) World Model Latency (ms) Policy Latency (ms)
[…]
Raw Sensor Stream Multi-view images 600 (BEVFormer) 2 2
Appendix H Visualization
H.1 Decoder Output by the World Model